OpenStudy (anonymous):

A force of 6N is exerted on a cart that is carrying a rock and produces an acceleration of 2 m/s^2. If the cart has a mass of 1 kg, what is the mass of the rock?

OpenStudy (anonymous):

F = m *a 6 = ( 1 + m) * 2 3 = 1 + m m = 2 kg

OpenStudy (anonymous):

Oh okay. Thanks :)

OpenStudy (anonymous):

Welcome...